# Efficient and Flexible Multirate Temporal Adaptivity

### v1

@@ 4 Fast Temporal Error Estimation
Since the $H\text{-}Tol$ multirate controller leverages an adaptive fast solver, we leverage the fact that at each substep of the fast integrator $t_{n,m}$, it computes an estimate of the local temporal error, $\varepsilon^{\mathfrak{f}}_{n,m}$ (e.g., using its own time step solution and embedding). Since the fast solver may be run with different tolerances than the slow solver, we denote its relative and absolute tolerances as $\text{reltol}^{\mathfrak{f}}_{n}$ and $\text{abstol}^{\mathfrak{f}}$, respectively. We then accumulate the local error estimates from each substep by assuming that none of these local errors cancel, estimating the total fast error as the sum of the substep errors,
@@
where $M$ is the set of all steps since the accumulator was last reset (generally at the start of the step, $t_{n}$). To understand the tolerance ratio, we note that both the fast and slow solvers use a weighted root-mean squared norm, such that an error norm of 1 corresponds with “sufficiently accurate,” i.e.,
$\|\varepsilon^{\mathfrak{f}}_{n,m}\|=\left(\frac{1}{N}\sum_{i=1}^{N}\left(\frac{\varepsilon^{\mathfrak{f}}_{n,m,i}}{\text{abstol}^{\mathfrak{f}}+\text{reltol}^{\mathfrak{f}}_{n}|y_{n,m,i}|}\right)^{2}\right)^{1/2},$
where $y_{n,m}$ is the time step solution at the start of the substep, and the vectors $y_{n,m}$ and $\varepsilon^{\mathfrak{f}}_{n,m}$ have $N$ components. The prefactor $\text{reltol}^{\mathfrak{f}}_{n}$ serves to convert the accumulated fast errors to a raw estimate of the relative error, before the factor $1/\text{reltol}^{\mathfrak{s}}_{n}$ scales this back to normalized relative error.
Since the $H\text{-}h$ multirate controller uses fixed fast substeps we cannot assume that the fast solver is adaptive. Thus, we use a more costly approach that runs the fast integrator using fixed time steps of size $h$ and $kh$ (e.g., with $k=2$) to achieve fast solutions $y_{n,h}^{\mathfrak{f}}$ and $y_{n,kh}^{\mathfrak{f}}$, which we then subtract to estimate the fast temporal error,
$\varepsilon^{\mathfrak{f},dbl}_{n}=\left|\frac{1}{k^{p}-1}\right|\;\|y_{n,m,h}^{\mathfrak{f}}-y_{n,m,kh}^{\mathfrak{f}}\|,$ (8)
where $p$ is the global order of accuracy for the method.
@@ 5 Numerical Tests
We use two benchmark problems. The first is a multirate, nonlinear version of the Kværno-Prothero-Robinson (KPR) ODE test problem, that has been slightly modified from [6, Section 6.2]:
@@
over $0<t<5$, where $p(t)=\cos(t)$ and $q(t)=\cos(\omega t(1+e^{-(t-2)^{2}}))$. This problem has analytical solution $u(t)=\sqrt{2+p(t)}$ and $v(t)=\sqrt{2+q(t)}$, and its behavior is dictated by the parameters: $e_{s}$ determines the strength of coupling from the fast to the slow scale, $e_{f}$ determines the coupling strength from the slow to the fast scale, $G<0$ determines the stiffness at the slow time scale, and $w$ determines the time-scale separation factor. We split the right-hand side above into up to three functions for the slow-implicit ($f^{\mathfrak{i}}$), slow-explicit ($f^{\mathfrak{e}}$) and fast ($f^{\mathfrak{f}}$) operators, respectively,
@@
for non-ImEx MRI methods the combined slow operator is $f^{\mathfrak{s}}=f^{\mathfrak{i}}+f^{\mathfrak{e}}$.
@@ 5.1 Embedded MRI Methods
The embeddings for each of the above methods have an order of accuracy one lower than the method itself. We note that the original MERK2, MERK3, MERK4, and MERK5 methods from [4] do not include embeddings. We provide embeddings for each in Appendix A.
@@ 5.2.1 Controller robustness
The $H\text{-}h$ controller family did not fare as well. Although some methods and controllers were able to meet the requested tolerances, most had errors that were orders of magnitude larger than requested. Generally, the $H\text{-}h$ controllers deteriorated as the problems became more multirate; in separate tests on weakly multirate problems (e.g., KPR with $\omega=5$ and the stiff Brusselator with $\epsilon=0.01$) this family was generally able to meet the requested error tolerances, which agrees with the results from [17]. The only outliers from this generally poor performance were the MERK21 and MERK32 methods, that achieved the target accuracy for all tolerances and multirate parameters. No other combination of adaptive MRI method and $H\text{-}h$ controller was able to achieve even within 10x of the target accuracy for the challenging problems with $\omega=500$ and $\epsilon=10^{-5}$. We cannot yet explain why the embedded MERK methods with $H\text{-}h$ controllers outperform other embedded MRI methods; however, we generally conclude that controllers in the $H\text{-}h$ family are considerably less robust than either the Decoupled or $H\text{-}Tol$ families for applications with nontrivial multirate nature, so we recommend that practitioners apply caution when using the $H\text{-}h$ family.
@@ 5.2.2 Adaptive MRI efficiency
MRI method performance. To compare individual embedded MRI methods, we downselected to the twelve fastest MRI method and controller pairs for each combination of problem, method order, and work metric. This process yielded eight sets of top-performing pairs. We then computed the intersection of the two sets corresponding to the different multirate values, retaining only those pairs that consistently rank among the top 12 for both. In Figures 2-4, we overlay the efficiency plots for these sets of “fastest” methods to determine the most performant MRI method and adaptivity controller pairs. In each figure, the KPR test problem is on the top row, and the stiff Brusselator test problem is on the bottom rows. Within each row, we show the plots for slow computational efficiency on the left.
@@ 6 Conclusions
In this work we present two new families of multirate time step adaptivity controllers, Decoupled and $H\text{-}Tol$, that are designed to work with embedded MRI methods for adapting time steps when solving problems with multiple time scales. Comparing against the previously-introduced $H\text{-}h$ controllers from [17], the proposed controllers offer dramatically improved performance and flexibility. While the proposed controller families are able to track multiple time scales for methods of a variety of orders of accuracy, the $H\text{-}h$ controllers seem to struggle to select steps with sufficient scale separation between $H$ and $h$, resulting in worse accuracy and higher cost than the proposed families. The combination of embedded MRI methods with the Decoupled and $H\text{-}Tol$ controllers theoretically support adaptive simulations of problems with an arbitrary number of time scales (here shown for a 3-scale benchmark problem), and achieve high accuracy while maintaining low computational cost. Of the proposed families, the $H\text{-}Tol$ controllers show much stronger performance for problems where cost is dominated by evaluation of the slow operators, at the expense of requiring estimates of the accumulated fast temporal error. On the other hand, the Decoupled controllers require no such accumulated fast error estimate, and show the best performance for problems where the fast and slow operators have comparable cost. Thus, we recommend that users consider this trade-off when selecting a multirate controller for their application.

### v2

@@ 4 Fast Temporal Error Estimation
Since the $H\text{-}Tol$ multirate controller relies on an adaptive fast solver, we leverage the fact that at each substep of the fast integrator $t_{n,m}$, it computes an estimate of the local temporal error, $\varepsilon^{\mathfrak{f}}_{n,m}$ (e.g., using its own time step solution and embedding). Since the fast solver may be run with a different relative tolerance than the slow solver, we denote this as $\text{reltol}^{\mathfrak{f}}_{n}$. As shown in (3.1), typical theory expects these local errors to grow exponentially based on the local Lipschitz constant, $L_{n}$; however in practice this accumulation is generally more well-behaved. In this paper, we choose a moderate accumulation assumption that the total fast error may be approximated as the sum of the substep errors,
@@
where $M$ is the set of all steps since the accumulator was last reset (generally at the start of the step, $t_{n}$). To understand the tolerance ratio, we assume that both the fast and slow solvers use a weighted root-mean squared norm (4), such that $\|\varepsilon^{\mathfrak{f}}_{n,m}\|\leq 1$ is “sufficiently accurate.” The prefactor $\text{reltol}^{\mathfrak{f}}_{n}$ serves to convert the accumulated fast errors to a raw estimate of the relative error, before the factor $1/\text{reltol}^{\mathfrak{s}}_{n}$ scales this back to normalized relative error.
We note that other accumulation strategies may be more well-suited to specific problems, where local errors may decay or even grow exponentially. In exploring options, we also considered a “maximum accumulation” strategy,
$\varepsilon^{\mathfrak{f},max}_{n}=\frac{\text{reltol}^{\mathfrak{f}}_{n}}{\text{reltol}^{\mathfrak{s}}_{n}}\max_{m\in M}\|\varepsilon^{\mathfrak{f}}_{n,m}\|,$ (17)
and an “average accumulation” strategy,
$\varepsilon^{\mathfrak{f},avg}_{n}=\frac{\text{reltol}^{\mathfrak{f}}_{n}}{\text{reltol}^{\mathfrak{s}}_{n}}\sum_{m\in M}\frac{h_{n,m}}{T_{M}}\|\varepsilon^{\mathfrak{f}}_{n,m}\|,$ (18)
where $T_{M}$ is the elapsed simulation time over the set $M$. In testing these approaches, we found that the additive strategy (16) showed the most robust performance; however, both of the strategies (17) and (18) are also included in the ARKODE time integration solver [22] from SUNDIALS [8, 9], where we have implemented our methods in this paper.
Since the $H\text{-}h$ multirate controller uses fixed fast substeps we cannot assume that the fast solver is adaptive, and thus it may not compute substep estimates $\|\varepsilon_{n,m}^{\mathfrak{f}}\|$. Thus, we use a more costly approach based on Richardson extrapolation to estimate the accumulated fast error over each slow MRI stage $i$. This runs the fast integrator using fixed time steps of size $h$ and $kh$ for $k\neq 1$, (e.g., $k=2$) to achieve fast solutions $v_{i,h}(\theta_{F,i})$ and $v_{i,kh}(\theta_{F,i})$, which we then subtract to estimate the fast temporal error over that slow stage,
$\varepsilon^{\mathfrak{f},dbl}_{i}=\left|\frac{1}{k^{p}-1}\right|\;\|v_{i,h}(\theta_{F,i})-v_{i,kh}(\theta_{F,i})\|,$
where $p$ is the global order of accuracy for the inner IVP solver [7, Sec. II.4]. We then take the maximum of these to estimate the overall fast temporal error,
$\varepsilon^{\mathfrak{f},dbl}_{n}=\max_{1\leq i\leq s}\varepsilon^{\mathfrak{f},dbl}_{i}.$ (19)
We note that this estimate may struggle for stiff problems, where order reduction may render the effective order to differ from $p$.
@@ 5 Numerical Tests
In our experience, MRI methods are most often used for two types of applications. The first are problems with multiple dynamcially-evolving time scales that must be resolved accurately, and where the fast time scale evolves considerably more rapidly than the slow time scale [24]. The second are problems with a stiff component, and where instead of applying an implicit or implicit-explicit (ImEx) solver to that term, it is subcycled with an explicit method [25]. We note that both [24] and [25] use MRI methods from ARKODE, but use fixed slow step sizes $H_{n}$ that could be improved based on the adaptive methods presented here.
To explore the performance of our adaptive methods, we test them on two benchmark problems that exhibit the aforementioned characteristics. The first is a multirate, nonlinear version of the Kværno-Prothero-Robinson (KPR) ODE test problem, that has been slightly modified from [6, Section 6.2]:
@@
over $0<t<5$, where $p(t)=\cos(t)$ and $q(t)=\cos(\omega t(1+e^{-(t-2)^{2}}))$. This problem has analytical solution $u(t)=\sqrt{2+p(t)}$ and $v(t)=\sqrt{2+q(t)}$, and its behavior is dictated by the parameters: $e_{s}$ determines the strength of coupling from the fast to the slow scale, $e_{f}$ determines the coupling strength from the slow to the fast scale, $G<0$ determines the stiffness at the slow time scale, and $w$ determines the time-scale separation factor. An illustration of the analytical solutions for this problem using parameters G=10G=-10, $e_{s}=e_{f}=1/10$, and a small time scale separation factor of $\omega=5$, is given in Figure 1 (left). We split the right-hand side above into up to three functions for the slow-implicit ($f^{\mathfrak{i}}$), slow-explicit ($f^{\mathfrak{e}}$) and fast ($f^{\mathfrak{f}}$) operators, respectively,
@@
for non-ImEx MRI methods the combined slow operator is $f^{\mathfrak{s}}=f^{\mathfrak{i}}+f^{\mathfrak{e}}$. We verify that the splitting (21) results in the desired time scales by running single scale simulations with $f^{\mathfrak{s}}$ and $f^{\mathfrak{f}}$ separately, with initial conditions $u(0)=v(0)=\sqrt{3}$ (the analytical solution values at $t=0$). We adaptively integrate each problem using MATLAB’s ode23 method, with absolute tolerance $10^{-11}$ and relative tolerance $10^{-6}$. In Figure 1 (right), we overlay the internal time steps taken to solve both problems, along with the internal steps required to solve the full problem, where we see that as expected these steps must adapt throughout each run, and also that the fast and slow scales are generally separated by a factor of $\omega$.
Figure 1: Left: example solution to the 2-scale KPR problem (20) with scale separation factor $\omega=5$. The oscillations in $v$ change frequency through the time interval, evolving approximately $\omega$ times faster than $u$. Right: internal time step sizes used when solving single-scale problems and full problem, showing the need for smaller step sizes when including the faster time scale.
@@
Figure 2 provides an illustration of the solutions for this problem using $\epsilon=1/2500$, as well as the time steps taken when solving the slow-only, fast-only, and full problems, using an identical approach as performed for Figure 1.
Figure 2: Left: example solution to the stiff Brusselator problem (22) with stiffness factor $\epsilon=1/2500$. Right: internal time step sizes used when solving single-scale problems and full problem. Since the fast time steps are primarily stability-limited, they remain relatively constant, while the slow time steps must rapidly adapt at first, and then remain steady for the remainder of the time interval.
@@ 5.1 Embedded MRI Methods
The embeddings for each of the above methods have an order of accuracy one lower than the method itself. We note that the original MERK2, MERK3, MERK4, and MERK5 methods from [4] do not include embeddings. We provide embeddings for each in A.
@@ 5.2.1 Controller robustness
The $H\text{-}h$ controller family did not fare as well. Although some methods and controllers were able to meet the requested tolerances, most had errors that were orders of magnitude larger than requested. Generally, the $H\text{-}h$ controllers deteriorated as the problems became more multirate; in separate tests on weakly multirate problems (e.g., KPR with $\omega=5$ and the stiff Brusselator with $\epsilon=0.01$) this family was generally able to meet the requested error tolerances, which agrees with the results from [17]. The only outliers from this generally poor performance were the MERK21 and MERK32 methods, that achieved the target accuracy for all tolerances and multirate parameters. No other combination of adaptive MRI method and $H\text{-}h$ controller was able to achieve even within 10x of the target accuracy for the challenging problems with $\omega=500$ and $\epsilon=10^{-5}$. We cannot yet explain why the embedded MERK methods with $H\text{-}h$ controllers outperform other embedded MRI methods; however, these results reinforce our analysis from Section 3.1.1 showing that the $H\text{-}h$ controllers struggle in comparison with the Decoupled and $H\text{-}Tol$ families for applications with significant multirate nature.
@@ 5.2.2 Adaptive MRI efficiency
MRI method performance. To compare individual embedded MRI methods, we downselected to the twelve fastest MRI method and controller pairs for each combination of problem, method order, and work metric. This process yielded eight sets of top-performing pairs. We then computed the intersection of the two sets corresponding to the different multirate values, retaining only those pairs that consistently rank among the top 12 for both. In Figures 4-6, we overlay the efficiency plots for these sets of “fastest” methods to determine the most performant MRI method and adaptivity controller pairs. In each figure, the KPR test problem is on the top row, and the stiff Brusselator test problem is on the bottom rows. Within each row, we show the plots for slow computational efficiency on the left.
@@ 6 Conclusions
In this work we present two new families of multirate time step adaptivity controllers, Decoupled and $H\text{-}Tol$, that are designed to work with embedded MRI methods for adapting time steps when solving problems with multiple time scales. Comparing against the previously-introduced $H\text{-}h$ controllers from [17], the proposed controllers offer dramatically improved performance and flexibility. While the proposed controller families are able to track multiple time scales for methods of a variety of orders of accuracy, the $H\text{-}h$ controllers struggle to select steps with sufficient scale separation between $H$ and $h$, resulting in worse accuracy and higher cost than the proposed families. The combination of embedded MRI methods with the Decoupled and $H\text{-}Tol$ controllers theoretically support adaptive simulations of problems with an arbitrary number of time scales (here shown for a 3-scale benchmark problem), and achieve high accuracy while maintaining low computational cost. Of the proposed families, the $H\text{-}Tol$ controllers show much stronger performance for problems where cost is dominated by evaluation of the slow operators, at the expense of requiring estimates of the accumulated fast temporal error. On the other hand, the Decoupled controllers require no such accumulated fast error estimate, and show the best performance for problems where the fast and slow operators have comparable cost. Thus, we recommend that readers consider this trade-off when selecting a multirate controller for their application.
@@
Additionally, the analysis in Section 3.1.1 hints at additional criteria that may be used to improve the $H\text{-}h$ controller family itself. Future studies can leverage this criteria to adjust the internal parameters of each $H\text{-}h$ controller to ensure their appropriateness for applications with large multirate ratios.